OpenStudy (anonymous):

how do you solve the system of equations 2x+y-3z=-3 x-2y+z=6 x-3y+3z=10

OpenStudy (amistre64):

either with substitution , or elimination, or matrix methods ...

OpenStudy (amistre64):

or, you could define the line of intersection of the first 2 planes and determine where it intersects the plane of the last one

OpenStudy (anonymous):

I came up with the answers x=34 y=-14 z=0 they worked in the first equation but not in the other two equations.

OpenStudy (amistre64):

then thats not a solution :)

OpenStudy (anonymous):

I don't know if I am doing something wrong when I try to calculate the other two equations because it did work in one of them x-2y+z=6 works

OpenStudy (amistre64):

when x=0 we have: 2y-6z=-6 -2y +z =6 ----------- 5z = 0; z=0, y=-3 (0,-3,0) ---------------------- when y=0 we have 2x-3z=-3 -2x-2z=-12 ------------ -5z = -15, z=3 , x=3 (3,0,3) ------------------------ the line between the points is therefore: x = 3+3t y = 0+3t z = 3+3t ----------------------- plug these into the last equation to determine the value of t x-3y+3z=10 (3+3t)-3(3t)+3(3+3t)=10 3+3t-9t+9+9t=10 3t=-2 , t = -2/3 ------------------- therefore x = 3+3(-2/3) y = 0+3(-2/3) z = 3+3(-2/3) x = 3-2 = 1 y = 0-2 = -2 z = 3-2 = 1 with any luck this worked oout without my typos

OpenStudy (amistre64):

2-2-3=-3 1+4+1=6 1+6+3=10 yep

OpenStudy (anonymous):

thank you so much

OpenStudy (amistre64):

i most likely used a method that you are not familiar with; but even so, the solution does not change :)